# Context and Diversity Matter: The Emergence of In-Context Learning in World Models

### v1

@@ Abstract
[…]
@@ 1 Introduction
The ability to predict future environmental states is crucial for reasoning and decision-making in both animals and humans. Inspired by this principle, constructing predictive models, especially the world model [1], to forecast environmental dynamics and outcomes forms the foundation for enabling agents to plan effective decisions and behaviors [2, 3, 4, 5, 6]. Consequently, world models are widely applied in fields such as navigation [7, 8, 9, 10, 11], autonomous driving [12, 13, 14, 15, 16], robotics [17, 18, 19, 20, 21, 22], and are considered a cornerstone of embodied artificial intelligence.
@@
Despite their proven effectiveness across various applications, previous prediction frameworks largely rely on static world models optimized for zero-shot, few-shot, or instantaneous performance. In contrast, humans and animals achieve real-time adaptation [23] through predictive coding—a process where prediction errors drive attention, generate feedback, and motivate learning and adjustment [24, 25]. For instance, when confronted with rare environments, humans experience surprise yet rapidly recalibrate their predictions for that setting, whereas static models continue to fail unless explicitly retrained on the relevant data. The ability to dynamically modify predictive mechanisms based on observational evidence, rather than relying solely on fixed parametric memory and external mapping modules, can effectively enable the model to adapt to environments unseen during training. This capability can be effectively addressed by In-Context Learning (ICL), as evidenced by recent advances in Large Language Models [26]. However, the potential of ICL for resolving plasticity in world models, particularly In-Context Environment Learning (ICEL), remains largely underexplored in current literature. Addressing this gap could further improve the generalization scope for embodied AI.
@@
We isolate two canonical mechanisms that let world models adapt in context. In-Context Environment Recognition (ICER) taps the model’s parametric memory of previously seen environments and uses the context only to identify which one is active. ICEL ignores prior settings and instead extracts the full environment from the context itself. By deriving error bounds for both modes, we show how their emergence depends on data distribution, including environment diversity, complexity, and trajectory length. To test the theory, we introduce Linear-attention Long-context world model, L2World, which effectively model long sequences with high computational efficiency. Across cart-pole control and vision-based indoor navigation, we find that distributional properties and long-context capacity govern whether the model invokes ICER or ICEL. Despite the light image encoders and decoders of L2World, it sets a new state-of-the-art for long-sequence observation prediction in navigation, outperforming approaches that rely on heavy diffusion-based image backbones. These results underscore the importance and the promise of deliberately incentivizing ICEL through carefully designed datasets and model architectures to strengthen the asymptotic generalization capability of world models, rather than focusing solely on instant or zero-shot frame-level performance.
@@ 2.1 Dynamic Prediction and World Models
Dynamic models, also referred to as world models [1, 27], encompass probabilistic, physical, or generative frameworks that formalize an AI system’s environmental understanding [28, 29, 30, 31]. These models predict future states by leveraging historical observations and play a pivotal role in advancing reinforcement learning (RL) methodologies and related domains. Specifically, they constitute foundational components in model-based RL [32, 33], enable simulations that facilitate agent learning through virtual experiences (thereby reducing reliance on direct environmental interaction) [34, 2, 5, 35, 36], and serve as auxiliary tasks to augment model supervision [37, 38, 39].
World models demonstrate robustness in integrating multi-modal raw sensor data, including visual, textual, inertial, and tactile inputs. To mitigate challenges posed by high-dimensional sensory inputs, representation learning paradigms such as Generative Adversarial Networks (GANs) [40] and Variational Autoencoders (VAEs) [41] are widely utilized to compress raw data into compact latent spaces. Subsequent temporal modeling in these reduced dimensions is achieved through latent world model architectures [2, 42, 4, 5, 43, 39, 44, 45], which capture sequential dependencies and enable coherent long-term predictions.
In navigation, early systems relied on traditional, hand-crafted pipelines such as SLAM. Recent work replaces these modules with generative world models, including diffusion [7], VAE [9], and RL-enhanced variants [46, 10], which reconstruct dynamics or simulate semantics [47, 48]. However, most existing methods disregard continual adaptation, especially across episodes, leaving a persistent gap between zero-shot performance and lifelong operation.
@@ 2.2 In-Context Learning and Meta-Learning
The approach based on parametric memory, which predominantly relies on gradient-based optimization or In-Weight Learning (IWL), has faced criticism for its lack of plasticity and the associated challenges it poses in continual learning scenarios [49]. Conversely, ICL has emerged as a pivotal capability in large language models [26], facilitating generalization to novel tasks without the necessity of parameter fine-tuning. ICL leverages contextual memory for task solutions, rather than depending on parametric memory. The concept of ICL is not novel. Meta-learning [50, 51], which focuses on acquiring learning capabilities rather than mastering specific skills, utilizes well-curated environments or data instead of relying on large-scale, uncurated pre-training data. Nevertheless, the lack of well-structured, task-rich, and cost-efficient datasets continues to present a significant challenge.
ICL has been utilized to encode a diverse array of learning mechanisms, including language learning [52], regression [53], Bayesian inference [54], reinforcement learning [55, 56, 57], and world models [58, 59], highlighting its resemblance to biological plasticity [60]. Yet existing work concentrates on few-shot in-context adaptation, overlooking ICL’s potential as contexts grow indefinitely. Moreover, little insight has been offered into its inner process and mechanisms. Concurrent studies identify at least two ICL modes in LLMs: task recognition (TR) and task learning (TL) [61]. Key factors conducive to ICL emergence have also been investigated [62, 58], such as “burstiness” and diversity, and so is the relation between IWL and ICL [63], but theory in this area remains under-developed.
@@ 3 Methodologies
Here, $D$ represents a metric measuring the error between two distributions (lower values indicate better performance). Notably, the ICL of the world model fundamentally relies on cross-episode contexts rather than intra-episode state estimation. To distinguish these concepts clearly, we use $q_{t}$ to denote short-term state estimation and $C_{T}$ to represent long-term ICL. While these are typically aligned in a single sequence in practice, maintaining this distinction facilitates rigorous theoretical analysis. Building on prior work that partitions in-context learning into task-recognition (TR) and task-learning (TL) modes [61, 64], theoretically, we are able to identify two analogous modes within world-model ICL: Environment Recognition (ER) and Environment Learning (EL). We then derive error bounds that characterize the conditions under which each mode emerges.
@@ 3.1 Environment Recognition (ER)
$\hat{p}_{\theta,ER}(o_{t+1}|q_{t},C_{T})$ $=\sum_{e\in\mathcal{E}}\underbrace{\hat{p}_{\theta}(e|q_{t},C_{T})}_{\text{Environment Recognition}}\cdot\underbrace{\hat{p}_{\theta,e}(o_{t+1}|q_{t})}_{\text{Environment-Specific World Model}}$ (3)
$\hat{p}_{\theta,e}(o_{t+1}|q_{t})$ $=\int\underbrace{\hat{p}_{\theta,s,e}(s_{t}|q_{t})}_{\text{State Estimation}}\cdot\underbrace{\hat{p}_{\theta,\tau,e}(s_{t+1}|s_{t},a_{t})}_{\text{Dynamics}}\cdot\underbrace{\hat{p}_{\theta,z,e}(o_{t+1}|s_{t+1})}_{\text{Observation Model}}\,ds_{t}$
Equation˜3 implies that in-context learning in world models arises mainly from the continual refinement of environment recognition, since the context yields no improvement to the environment-specific world-model term $\hat{p}_{\theta,e}=\{\hat{p}_{\theta,s,e},\hat{p}_{\theta,\tau,e},\hat{p}_{\theta,z,e}\}$. Therefore, in the ER regime, the model first acquires world models for the entire environment set through IWL or parametric memory, and then uses the context solely to identify the current environment.
@@ 3.2 Environment Learning (EL)
An intuitive observation of Equation˜4 is that EL functions, at minimum, as a in-context memorizer. For highly complex environments, its performance might degrade sharply, because accurate estimation of the transition of $q_{t}$ demands contexts that scale with the environment’s complexity.
@@ 3.3 Theoretical Analyses of ER and EL
Although the training process and data distribution play a key role in effectively incentivizing ICL [62], how does the data distribution determine whether EL or ER emerges? If training consistently minimizes predictive error, the error bounds of EL and ER become the decisive factor in selecting the emergent mode. To investigate the conditions governing the emergence of the two modes (ER and EL), we analyze the error upper bounds for each paradigm. For tractable analysis, we introduce the following simplifying assumptions: (1) The observation, state, and action spaces are discrete; (2) Both modes achieve ideal state estimation $p_{e}(s|q)$; (3) The context $C_{T}$ has a uniform state-action distribution. Under these assumptions, we derive an upper bound on the error of the world model optimized over environment set $\mathcal{E}$, measured by the total variation (TV) distance, when deployed in an unseen environment $e_{0}$ at context horizon $T$. Formally, the TV error is bounded by:
@@
$\text{TV}(\hat{p}_{ER},p_{e_{0}})\leq$ $\min\{\underbrace{\alpha(|\mathcal{E}|-1)T^{-1/2}/3}_{\text{Identification Error}}+\underbrace{\min_{e\in\mathcal{E}}TV(\hat{p}_{\theta,e},p_{e_{0}})}_{\text{Best Matching Error}},\quad\underbrace{\max_{e\in\mathcal{E}}TV(\hat{p}_{\theta,e},p_{e_{0}})}_{\text{Worst Matching Error}}\}$
@@
Proofs and detailed assumptions for the above theorems are deferred to Appendix˜A. An immediate observation from Theorem˜1 is that ICEL enjoys an ideal error upper bound that decays as $T^{-1/2}$, whereas ICER carries a non-decaying residual term (the best-matching error) that becomes the dominant obstacle to generalizing across unseen environments. To enhance generalization, we therefore ask: under what condition is ICEL preferred to ICER? For the entire training set $\mathcal{E}$, ICEL dominates whenever $\mathbb{E}_{e\in\mathcal{E}}\bigl[\text{TV}(\hat{p}_{\text{EL}},p_{e})\bigr]\ll\mathbb{E}_{e\in\mathcal{E}}\bigl[\text{TV}(\hat{p}_{\text{ER}},p_{e})\bigr];$ The opposite inequality favors ICER. Although the errors themselves are intractable to evaluate directly, the following insights are obtained by comparing their upper bounds:
(1) Lower environment complexity and greater environment diversity favor EL over ER: Note that the best-matching error is effectively zero, because the model is always evaluated on environments seen during training. The cardinality of the training set, $|\mathcal{E}|$, enters only the ER bound, whereas the environment complexity, $|O||S||A|$, enters only the EL bound. Consequently, lower complexity and a larger training set push the EL bound below the ER bound. Moreover, because ER error is limited by the worst-matching error, scarce diversity in the training environments can keep the ER error small and prevent EL from ever overtaking it.
(2) Long context and environment diversity are key to both ER and EL: Insufficient diversity not only blocks the emergence of ICEL but also suppresses ICER, because a small worst-matching error removes the pressure to reduce the identification error. Once the training set is sufficiently diverse, both ER and EL obtain an upper error bound that decays as $T^{-1/2}$, demonstrating that long context is indispensable for either mechanism.
(3) Over-training and powerful IWL facilitate ER over EL: we hypothesize that IWL perfectly models the environment-specific dynamics ($\hat{p}_{\theta,e}$) in the training set, so the best-matching error is nearly zero during training; however, this is not always true. Early in training, IWL can still incur large errors, transiently pushing the model toward EL instead of ER. This transiency is also investigated by prior works in ICL [65]. As training proceeds and IWL becomes increasingly accurate, the model may revert to ER.
In the following section, we empirically confirm that these insights hold not only for discrete settings but also for continuous MDPs and POMDPs. Because theory predicts that large environmental diversity and low task complexity are required for incentivizing EL, and most of the closed benchmarks can not satisfy those requirements. We construct our dataset from randomly sampled cart-poles and procedurally generated mazes to evaluate the performance of ICER and ICEL.

### v2

# arXiv:2509.22353v2[cs.LG] 27 Feb 2026

@@ Abstract
The capability of predicting environmental dynamics underpins both biological neural systems and general embodied AI in adapting to their surroundings. Yet prevailing approaches rest on static world models that falter when confronted with novel or rare configurations. We investigate in-context learning (ICL) of world models, shifting attention from zero-shot performance to the growth and asymptotic limits of the world model. Our contributions are three-fold: (1) we formalize ICL of a world model and identify two core mechanisms: environment recognition (ER) and environment learning (EL); (2) we derive error upper-bounds for both mechanisms that expose how the mechanisms emerge; and (3) we empirically confirm that distinct ICL mechanisms exist in the world model, and we further investigate how data distribution and model architecture affect ICL in a manner consistent with theory. These findings demonstrate the potential of self-adapting world models and highlight the key factors behind the emergence of EL/ER, most notably the necessity of long context and diverse environments. The codes are available at https://github.com/airs-cuhk/airsoul/tree/main/projects/MazeWorld.
@@ 1 Introduction
The ability to predict future environmental states is crucial for reasoning and decision-making in both animals and humans. Inspired by this principle, constructing predictive models, especially the world model Ha and Schmidhuber (2018a), to forecast environmental dynamics and outcomes forms the foundation for enabling agents to plan effective decisions and behaviors Hafner et al. (2025); Zhang et al. (2023a); Mazzaglia et al. (2024); Samsami et al. (2024); Alonso et al. (2024). Consequently, world models are widely applied in fields such as navigation Bar et al. (2025); Mendonca et al. (2021); Koh et al. (2021); Duan et al. (2024); Liu et al. (2025), autonomous driving Hu et al. (2023); Russell et al. (2025); Gao et al. (2024); Zhang et al. (2024a); Wang et al. (2024c), robotics Zhang et al. (2023b); Wu et al. (2023); Hansen et al. (2024); Pang et al. (2025); Zhou et al. (2024); Barcellona et al. (2025), and are considered a cornerstone of embodied artificial intelligence.
@@
Despite their proven effectiveness across various applications, previous prediction frameworks largely rely on static world models optimized for zero-shot, few-shot, or instantaneous performance. In contrast, humans and animals achieve real-time adaptation Vorhees and Williams (2014) through predictive coding—a process where prediction errors drive attention, generate feedback, and motivate learning and adjustment Rao and Ballard (1999); Salvatori et al. (2023). For instance, when confronted with rare environments, humans experience surprise yet rapidly recalibrate their predictions for that setting, whereas static models continue to fail unless explicitly retrained on the relevant data. The ability to dynamically modify predictive mechanisms based on observational evidence, rather than relying solely on fixed parametric memory and external mapping modules, can effectively enable the model to adapt to environments unseen during training. This capability can be effectively addressed by In-Context Learning (ICL), as evidenced by recent advances in Large Language Models (LLMs) Brown et al. (2020). However, existing investigations into ICL primarily focus on language-related tasks or simpler few-shot classification and regression tasks. ICL within world models, despite its potential importance, remains largely underexplored. Addressing this gap could complete the missing pieces of embodied AI, further improving generalization and self-adaptivity.
@@
Following the Bayesian hypothesis of ICL Panwar et al. (2023); Xie et al. (2024), we clarify two potential underlying mechanisms for ICL in world models: environment recognition (ER), which relies on parametric memory of the training environment, and environment learning (EL), which does not. By deriving upper error bounds for both ICL modes, we theoretically demonstrate that the emergence of ICL depends on environment diversity, complexity, and context length. This insight motivates the development of long-context adaptive world modeling. Consequently, we introduce the linear-attention long-context world model, L2World, which enables self-adaptation to environments through efficient memory updates within the context. Across cart-pole control and vision-based indoor navigation tasks, we empirically demonstrate that distributional properties and long-context capacity govern the ICL ability of world models. Despite employing lightweight image encoders and decoders in L2World, it establishes a new state-of-the-art for long-sequence observation prediction in cross-environment adaptation, outperforming methods that rely on computationally intensive diffusion-based image backbones. These results underscore the importance and potential of enhancing ICL through intentionally diversified datasets and long-context modeling architectures within world models, rather than focusing solely on immediate or zero-shot frame-level performance.
@@ 2.1 Dynamic Prediction and World Models
Dynamic models, also referred to as world models (Ha and Schmidhuber, 2018a; Forrester, 1958), encompass probabilistic, physical, or generative frameworks that formalize an AI system’s environmental understanding (Sutton, 1990; Battaglia et al., 2013; Ha and Schmidhuber, 2018b; LeCun, 2022). These models predict future states by leveraging historical observations and play a pivotal role in advancing reinforcement learning (RL) methodologies and related domains. Specifically, they constitute foundational components in model-based RL (Finn and Levine, 2017; Schrittwieser et al., 2020), enable simulations that facilitate agent learning through virtual experiences (thereby reducing reliance on direct environmental interaction) (Hafner et al., 2019a; 2025; Samsami et al., 2024; Sutton, 1991; Kaiser et al., 2020), and serve as auxiliary tasks to augment model supervision (Jaderberg et al., 2016; Hu et al., 2022; Zhang et al., 2024c).
World models demonstrate robustness in integrating multi-modal raw sensor data, including visual, textual, inertial, and tactile inputs. To mitigate challenges posed by high-dimensional sensory inputs, representation learning paradigms such as Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) and Variational Autoencoders (VAEs) (Kingma and Welling, 2013) are widely utilized to compress raw data into compact latent spaces. Subsequent temporal modeling in these reduced dimensions is achieved through latent world model architectures (Hafner et al., 2025; Wang et al., 2024d; Mazzaglia et al., 2024; Samsami et al., 2024; Hafner et al., 2019b; Zhang et al., 2024c; Garrido et al., 2024; Li et al., 2024), which capture sequential dependencies and enable coherent long-term predictions.
In navigation, early systems relied on traditional, hand-crafted pipelines such as SLAM. Recent work replaces these modules with generative world models, including diffusion Bar et al. (2025), VAE Koh et al. (2021), and RL-enhanced variants Poudel et al. (2023); Duan et al. (2024), which reconstruct dynamics or simulate semantics Nie et al. (2025); Liu et al. (2023). However, most existing methods disregard continual adaptation, especially across episodes, leaving a persistent gap between zero-shot performance and lifelong operation.
@@ 2.2 In-Context Learning and Meta-Learning
The approach based on parametric memory, which predominantly relies on gradient-based optimization or In-Weight Learning (IWL), has faced criticism for its lack of plasticity and the associated challenges it poses in continual learning scenarios Dohare et al. (2024). Conversely, ICL has emerged as a pivotal capability in large language models Brown et al. (2020), facilitating generalization to novel tasks without the necessity of parameter fine-tuning. ICL leverages contextual memory for task solutions, rather than depending on parametric memory. The concept of ICL is not novel. Meta-learning Santoro et al. (2016); Duan et al. (2016), which focuses on acquiring learning capabilities rather than mastering specific skills, utilizes well-curated environments or data instead of relying on large-scale, uncurated pre-training data. Nevertheless, the lack of well-structured, task-rich, and cost-efficient datasets continues to present a significant challenge.
ICL has been utilized to encode a diverse array of learning mechanisms, including language learning Akyürek et al. (2024), regression Garg et al. (2022), reinforcement learning Laskin et al. (2023); Lee et al. (2023); Wang et al. (2025), and world models Anand et al. (2022); Gupta et al. (2024), highlighting its resemblance to biological plasticity Lior et al. (2024). Yet existing work concentrates on few-shot in-context adaptation, overlooking ICL’s potential as contexts grow indefinitely. Concurrent studies have identified various mechanisms and circuits underlying ICL, including distinctions between task learning and task recognition Pan et al. (2023), as well as retrieval versus inference Park et al. (2025), among others. Key factors influencing the emergence of ICL have also been investigated, such as transience, task diversity, and context length Chan et al. (2022); Anand et al. (2022); Wurgaft et al. (2025); Nguyen and Reddy (2025), along with the relationship between IWL and ICL Chan et al. (2025); Singh et al. (2025). However, these studies primarily focus on simplified regression and classification problems, while theoretical frameworks addressing the incentivization of ICL within world models remain underexplored.
@@ 3 Methodologies
Here, $D$ represents a metric measuring the error between two distributions (lower values indicate better performance). Notably, the ICL of the world model fundamentally relies on cross-episode contexts rather than intra-episode state estimation. To distinguish these concepts clearly, we use $q_{t}$ to denote short-term state estimation and $C_{T}$ to represent long-term ICL. While these are typically aligned in a single sequence in practice, maintaining this distinction facilitates rigorous theoretical analysis. Building on prior work that partitions in-context learning into different modes Kirsch et al. (2022); Pan et al. (2023); Park et al. (2025), theoretically, we are able to identify two analogous modes within world-model ICL: Environment Recognition (ER) and Environment Learning (EL). We then derive error bounds that characterize the conditions under which each mode emerges.
@@ 3.1 Environment Recognition (ER)
$\hat{p}_{\theta,ER}(o_{t+1}|q_{t},C_{T})=\sum_{e\in\mathcal{E}}\underbrace{\hat{p}_{\theta}(e|q_{t},C_{T})}_{\begin{subarray}{c}\text{Environment}\\ \text{Recognition}\end{subarray}}\cdot\underbrace{\hat{p}_{\theta,e}(o_{t+1}|q_{t})}_{\begin{subarray}{c}\text{Environment-Specific}\\ \text{World Model}\end{subarray}}$ (3)
Therefore the ICL in world models arises mainly from the recognition of seen environments, while environment-specific world model are static within the inference. The environment-specific world model can be further decomposed into:
$\hat{p}_{\theta,e}(o_{t+1}|q_{t})=\int\underbrace{\hat{p}_{\theta,s,e}(s_{t}|q_{t})}_{\text{State Estimation}}\cdot\underbrace{\hat{p}_{\theta,\tau,e}(s_{t+1}|s_{t},a_{t})}_{\text{Dynamics}}\cdot\underbrace{\hat{p}_{\theta,z,e}(o_{t+1}|s_{t+1})}_{\text{Observation Model}}\,ds_{t},$ (4)
where the estimate $\hat{p}_{\theta,e}$ is obtained by marginalizing over the terms that include state estimation, dynamics, and observation $({\hat{p}_{\theta,s,e},\hat{p}_{\theta,\tau,e},\hat{p}_{\theta,z,e}})$. Therefore, in the ER regime, the model first acquires world models for the entire environment set through IWL or parametric memory, and then uses the context solely to identify the current environment.
@@ 3.2 Environment Learning (EL)
An intuitive observation of Equation˜5 is that EL functions, at minimum, as an in-context memorizer. For highly complex environments, its performance might degrade sharply, because accurate estimation of the transition of $q_{t}$ demands contexts that scale with the environment’s complexity.
@@ 3.3 Theoretical Analyses of ER and EL
Although the training process and data distribution play a key role in effectively incentivizing ICL Chan et al. (2022), how does the data distribution determine whether EL or ER emerges? If training consistently minimizes predictive error, the error bounds of EL and ER become the decisive factor in selecting the emergent mode. To investigate the conditions governing the emergence of the two modes (ER and EL), we analyze the error upper bounds for each paradigm. For tractable analysis, we introduce the following simplifying assumptions: (1) The observation, state, and action spaces are discrete; (2) Both modes achieve ideal state estimation $p_{e}(s|q)$; (3) The context $C_{T}$ has a uniform state-action distribution. Under these assumptions, we derive an upper bound on the error of the world model optimized over environment set $\mathcal{E}$, measured by the total variation (TV) distance, when deployed in an unseen environment $e_{0}$ at context horizon $T$. Formally, the TV error is bounded by:
@@
$\text{TV}(\hat{p}_{ER},p_{e_{0}})\leq$ $\min[\underbrace{\alpha/3\cdot(|\mathcal{E}|-1)\cdot T^{-1/2}}_{\text{Recognition Error}},\underbrace{\max_{e_{1},e_{2}\in\mathcal{E}}TV(p_{e_{1}},p_{e_{2}})}_{Diversity}]+\underbrace{\min_{e\in\mathcal{E}}TV(\hat{p}_{\theta,e},p_{e_{0}})}_{\text{Best Matching Error}}$
@@
Proofs and detailed assumptions for the above theorems are deferred to Appendix˜A. An immediate observation from Theorem˜1 is that EL enjoys an ideal error upper bound that decays as $T^{-1/2}$, whereas ER carries a non-decaying residual term (the best-matching error) that becomes the dominant obstacle to generalizing across unseen environments. To enhance generalization, we therefore ask: under what condition is EL preferred to ER? For the entire training set $\mathcal{E}$, EL dominates whenever $\mathbb{E}_{e\in\mathcal{E}}\bigl[\text{TV}(\hat{p}_{\text{EL}},p_{e})\bigr]\ll\mathbb{E}_{e\in\mathcal{E}}\bigl[\text{TV}(\hat{p}_{\text{ER}},p_{e})\bigr];$ The opposite inequality favors ER. Although the errors themselves are intractable to evaluate directly, the following insights are obtained by comparing their upper bounds:
(1) Lower environmental complexity and a greater number of environments favor EL over ER: Note that the best-matching error is effectively zero because the model is evaluated only on environments seen during training. The cardinality of the training set, $|\mathcal{E}|$, affects only the ER bound, whereas the environmental complexity, $|O||S||A|$, influences only the EL bound. Consequently, lower complexity combined with a larger training set pushes the EL bound below the ER bound.
(2) Long context and environment diversity are key to both ER and EL: As the upper error bound of ER effectively approaches zero when diversity is low, the emergence of both ER (where the identification error would never dominate) and EL is precluded. Once the training set is sufficiently diverse, both ER and EL obtain an upper error bound that decays as $T^{-1/2}$, demonstrating that long context is indispensable for either mechanism.
(3) Over-training and powerful IWL facilitate ER over EL: we hypothesize that IWL perfectly models the environment-specific dynamics ($\hat{p}_{\theta,e}$) in the training set, so the best-matching error is nearly zero during training; however, this is not always true. Early in training, IWL can still incur large errors, transiently pushing the model toward EL instead of ER. This transiency is also investigated by prior works in ICL Singh et al. (2023; 2025). As training proceeds and IWL becomes increasingly accurate, the model may revert to ER.
In the following section, we empirically confirm that these insights hold not only for discrete settings but also for continuous MDPs and POMDPs. Because theory predicts that large environmental diversity and low task complexity are required for incentivizing EL, and most of the closed benchmarks can not satisfy those requirements. We construct our dataset from randomly sampled cart-poles and procedurally generated mazes to evaluate the performance of ER and EL.